How fast is 5L/min, the rate at which your heart pumps blood? Is it closer to a sink faucet, a water fountain or a bathtub spigot? Describe the method that you used to determine answer and your reasoning.
Answer #1

1L = 10^-3 m^3

5L = 5*10^-3 m^3

1 min = 60 s

= 5*10^-3 / 60

= 8.33*10^-5 m^3/s

so the discharge rate is 8.33*10^-5 m^3/s